OpenStudy (anonymous):

how does a person factor this ? 56xy - 60mx -24x^2 + 140my

OpenStudy (anonymous):

Very similar in a way the person who has made it....

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

Nothing..

OpenStudy (anonymous):

well you can start by factoring numbers out :P

OpenStudy (raden):

56xy - 60mx -24x^2 + 140my = 4(14xy - 15mx -6x^2 + 35my) = 4(14xy+35my - (15mx+35my)) = 4((7y)(2x+5m) - 3x(2x+5m)) = 4(2x+5m)(7y-3x)

OpenStudy (anonymous):

\[56xy - 60mx -24x^2 + 140my \implies 56xy + 140my - 24x^2 - 60mx\] \[7y(8x + 20m) -3 (8x + 20m) \implies (7y - 3)(8x + 20m)\] \[\implies \color{blue}{4(7y - 3)(2x + 5m)}\]

OpenStudy (anonymous):

thanks for your help, all of you. your examples are relieving. it all seems so simple now. :)